# ALIFE2024 template

### v2

@@ Abstract
A major challenge when describing the origin of life is to explain “how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013). So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined. Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication). In contrast to other theories that assume the existence of initial complex structures, our representation starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines. By describing precisely the primordial transitions in chemistry-based computation, our framework is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales. At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
@@ Introduction
The undecidability behind self-replication lead us explicitly to open-ended evolution as a result of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit (strong) open-ended evolution must be undecidable. The PDA allows the ecosystem to store a more complex set of instructions—including the components of the chemical network that makes up the PDA itself—in a delocalized way. Thus, in this evolutionary phase the system is capable of developing extra functions, which gives to the PDA the ability of self-preserving behavior and propagate mutations to its offspring. This is consistent with the fact that a simple catalytic reaction network may be able of encoding correlations in environmental variables, evolving some kind of associative learning (Bartlett and Louapre,, 2022). Following the law of requisite variety (Ashby,, 1956), these chemical networks evolve concurrently to the environment in which they were developed, producing more semantic information which translates as a primitive degree of agency (Kolchinsky and Wolpert,, 2018). In addition, observe that the transition between FSAs and one-stack PDAs is compatible with ideas related to the appearance composomes (Segré and Lancet,, 2000; Hunding et al.,, 2006) and their maintenance (Markovitch and Lancet,, 2012).
@@ Acknowledgements
The authors acknowledge Luis M. Rocha and three anonymous reviewers for their valuable comments during the review of this manuscript.

### v3

@@ Abstract
A major challenge when describing the origin of life is to explain “how instructional information control systems emerge naturally and spontaneously from mere molecular dynamics” (Walker and Davies,, 2013). So far, no one has clarified how information control emerged ab initio and how primitive control mechanisms in life might have evolved, becoming increasingly refined. Based on recent experimental results showing that chemical computation does not require the presence of life-related chemistry (Dueñas-Díez and Pérez-Mercader,, 2019), we elucidate the origin and early evolution of information handling by chemical automata, from information processing (computation) to information storage (memory) and information transmission (communication) and later digital messengers, covering at the same time its syntactic, semantic and pragmatic flavors (Küppers,, 1990). In contrast to other theories that assume the existence of initial complex structures, our representation starts from trivial self-replicators whose interaction leads to the arising of more powerful molecular machines. By describing precisely the primordial transitions in chemistry-based computation, our framework is capable of explaining the above-mentioned gaps and can be translated to other models of computation, which allow us to explore biological phenomena at multiple spatial and temporal scales. Being compatible with the free energy principle, we have developed a computational enactivist (Korbak,, 2021) theoretical framework that could be able to describe from the origin of life to higher-level cognition, as if it were a purely constructivist narrative (Stewart,, 1995). At the end of our manuscript, we propose some ways to extend our ideas, including experimental validation of our theory (both in vitro and in silico).
@@ Introduction
The undecidability behind self-replication lead us explicitly to open-ended evolution (OEE) as a result of evolution (Pattee and Sayama,, 2019), and as Hernández-Orozco et al., (2018) has pointed out, systems that exhibit strong OEE must be undecidable. The PDA allows the ecosystem to store a more complex set of instructions—including the components of the chemical network that makes up the PDA itself—in a delocalized way. Thus, in this evolutionary phase the system is capable of developing extra functions, which gives to the PDA the ability of self-preserving behavior and propagate mutations to its offspring. This is consistent with the fact that a simple catalytic reaction network may be able of encoding correlations in environmental variables, evolving some kind of associative learning (Bartlett and Louapre,, 2022). Following the law of requisite variety (Ashby,, 1956), these chemical networks evolve concurrently to the environment in which they were developed, producing more semantic information which translates as a primitive degree of agency (Kolchinsky and Wolpert,, 2018). In addition, observe that the transition between FSAs and one-stack PDAs is compatible with ideas related to the appearance composomes (Segré and Lancet,, 2000; Hunding et al.,, 2006) and their maintenance (Markovitch and Lancet,, 2012).
@@
We are converging towards a technological singularity in which the digital is dominant, in such a way that capital is decentralized due to the ease of societies to acquire material goods that allow them to access this digital world. In this way, we are heading towards a collective decentralization, in which even material goods will have to be decentralized in order to continue with the socio-techno-economic progress we have seen in the last decades, given by universal indicators such as the global increase in life expectancy, among others.
@@ Acknowledgements
The authors acknowledge Juan Pérez-Mercader, Luis M. Rocha and three anonymous reviewers for their valuable comments during the evaluation of this manuscript.